40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5769     Entry time: Mon Oct 31 14:01:56 2011
Author: Katrin 
Type: Update 
Category: Green Locking 
Subject: YARM locks for 2h 

Today, I could lock the YARM laser for 2h to the YARM cavity. After to hours the output of the servo is saturated. I need to work on thermal feedback to the laser.

It is a nice TEM00 mode and the green light enters PSL table.

20111031_OLTF.png

Measured with pin-ball machine spectrum analyzer (I forgot the real name, but it is the one that makes sounds like a pin-ball machine), source power10mVp, Lb1005 gain 2.05.

 

Setup

YARM_setup.png

Input offset of LB1005 is zero

 

Locking history

On Thursday, Oct 27, lock for 3 min

On Friday, Oct 28, lock up to 18 min, improvements done by

  • finding the right adjustment of PI Corner frequency and gain
  • better alignment of the light to the cavity
  • I used a high-pass filter between LO and LB1005, but no improvement of lock. In contrast, it got worse.

On Monday,Oct 31, careful adjustment of summing box (rear of of LB1005), lock up to 2h, limited by saturated feedback signal --> work on slow control

 

Some more plots

feedback.png

 PD_DC1.png

PDH_error1.png

Attachment 1: 20111031_OLTF.png  8 kB  | Show | Show all
Attachment 2: YARM_setup.png  649 kB  Uploaded Mon Oct 31 15:37:39 2011  | Show | Show all
ELOG V3.1.3-